OpenStudy (anonymous):

if y = x-3 , which ordered pair is NOT a solution to the equation y=x-3 a. (2,-1) b. (-3,-6) c. (3,0) d. (-4,-6)

OpenStudy (anonymous):

You simply plug in [x, y] values into the equation. For example, y = x - 3 a. (2, -1) -1 = 2 - 3 -1 = -1 So we know that a is an ordered pair. Now you try out for the rest.

OpenStudy (anonymous):

Choice (D) does not satisfy that equation becuase when x = -4 and y = -6, - 6 does not equal to -4 -3 -6 does not equal to - 7.

OpenStudy (anonymous):

Easyaspi314 has already given you the answer. Though, what good is the answer if you don't know how to do it yourself? You simply plug in values for x and y into the equation and if it does not equal, then it is not an ordered pair.

OpenStudy (anonymous):

is it b or d ?